OpenStudy (anonymous):

Plzz help !! ..Write the fraction or mixed number as a decimal. Use a bar to show a repeating decimal. 1 1 - 4

OpenStudy (anonymous):

b a - = c*a+b /c(In general ) c Here a=1,b=1,c=4 So this will give (4*1+1)/4 =5/4 =1.25 Ans

OpenStudy (anonymous):

thnxx